OpenStudy (anonymous):

The radius of one sphere is twice as great as the radius of a second sphere. The ratio of the volume of the smaller sphere to the volume of the larger sphere is _____.

OpenStudy (ganpat):

let radius of sphere 1 = x and for other will be 2x Their volume is given by 4/3 * Pi * r3.. So ratio of volume = V1/V2 = r1 cube/ r2 cube = x cube /(2x) cube = 1/8..